Step-by-Step Explanation for this problem please:
Show that the equations are equal by rewriting them in the form of A(t) = ab^t:
A(t) = 100e^(-0.005t) and A(t) = 100(1/2)^(t/138)

Answer by stanbon(75887)

Show that the equations are equal by rewriting them in the form of A(t) = ab^t:
A(t) = 100e^(-0.005t)
A(t) = 100e^(-0.005)t
A(t) = 100(0.995012)^t
===========================
A(t) = 100(1/2)^(t/138)
A(t) = 100(1/2)^(1/138)t
A(t) = 100(0.994990)^t
===========================
Cheers,
Stan H.